ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 173, 2009

Banks, W. D. and I. E. Shparlinski Sato–Tate, cyclicity, and divisibility statistics on average for elliptic curves of small height

253

Ben Yaacov, I. Continuous and random Vapnik–Chervonenkis classes

309

Boutet de Monvel, A., D. Lenz and P. Stollmann Sch’nol’s theorem for strongly local forms

189

Cohen, A. M., B. Frenk and D. B. Wales Brauer algebras of simply laced type

335

Collins, D. J. and A. Juhasz The geometry of the consequences of a single Engel relator

125

Dahmani, F. Existential questions in (relatively) hyperbolic groups  91

Frenk, B. See A. M. Cohen Hamada, H., T. Honda and G. Kohr Bohr’s theorem for holomor-

phic mappings with values in homogeneous balls

177

Honda, T. See H. Hamada Juhasz, A. See D. J. Collins del Junco, A. and A. Şahin Dye’s theorem in the almost continuous

category

235

Kochloukova, D. H. and A. G. S. Pinto Finiteness conditions on subgroups of profinite p-Poincaré duality groups

367

Kohr, G. See H. Hamada

Kojman, M. A symmetrized metric Ramsey theorem  305

Lenz, D. See A. Boutet de Monvel Morbidelli, D. Liouville theorem, conformally invariant cones and

umbilical surfaces for Grushin-type metrics

379

Pinto, A. G. S. See D. H. Kochloukova Rubin, B. The Lower Dimensional Busemann-Petty problem for bod-

ies with the generalized axial symmetry

213

Şahin, A. See A. del Junco

Samet, I. Rigid actions of amenable groups  61

Schippers, E. and W. Staubach Variation of Neumann and Green functions under homotopies of the boundary

279

Shelah, S. Dependent first order theories, continued  1

Shparlinski, I. E. See W. D. Banks Spurný, J. On the Dirichlet problem of extreme points for non-

continuous functions

403

Staubach, W. See E. Schippers Stollmann, P. See A. Boutet de Monvel Urban, R. and J. Zienkiewicz Dimension free estimates for Riesz

transforms of some Schrödinger operators

157

Wales, D. B. See A. M. Cohen Zienkiewicz, J. See R. Urban Zweimüller, R. Measure preserving transformations similar to

Markov shifts 421